### v1

@@ 4.3 The algorithmic construction
noticing that the running time of Theorem 4.3 (and thus of Theorems 2.14 and 2.15) is inversely proportional to $\varepsilon$, and the increase in the original dependency of $k$ needed for the construction is directly proportional to $\alpha$. Since $\alpha$ has to be at least $1$ for the choices of $b$, $x$, and $d$ above to be feasible for the construction, and $\varepsilon>0$ is needed to apply Theorem 4.3, we conclude that $10<\varepsilon<$ 1 and $\alpha>1$. Taking $\alpha$ and $\varepsilon$ satisfying those bounds and observing that $|\mathcal{I}|\geq 0.6|V|\geq 0.6a(a-1)$ (since in $V$ we excluded only the pairs $(i,i)$ with $i\in[a]$), we apply 4.6 and finish the construction.
In case 2., the goal is to apply 4.5 to a path system built in three steps. For each each $e\in M_{1}\cup M_{2}$ with endpoints $(i,j)$ and $(i^{\prime},j^{\prime})$, the Bowtie Lemma [19, Lemma 3.7] is applied to $L_{i,j}$ and $L_{i^{\prime},j^{\prime}}$ to, in short, construct sets of closed walks with bounded congestion assuming that the intersection graph of the two linkages has sufficiently large degeneracy. By the definitions of $H_{1}$ and $H_{2}$, for each $\ell\in\{1,2\}$ and $e\in M_{\ell}$ the intersection graph of the pair of threaded linkages associated with the endpoints of $e$ is not $d_{\ell}$-degenerate, each such edge is associated with a set $Z_{e}$ of closed walks of congestion at most $4$ with $|Z_{e}|\geq d_{\ell}/(2^{9}\cdot 5)$ (the denominator comes from the statement of the Bowtie Lemma). By construction of $Z$, the same holds for each $(i,j)\in Z$.
@@
If, for some $F\subseteq Z\cup M_{1}\cup M_{2}$ of size $1$ or $2$ the intersection graph ${\sf Int}(\bigcup_{g\in F}Z_{g})$ is not $d_{3}$-degenerate, then immediately apply 4.7 to construct the desired bramble and the algorithm stops. Thus, we now assume that for any such $F$ the graph ${\sf Int}(\bigcup_{g\in F}Z_{g})$is $d_{3}$-degenerate.
@@
Similarly to case 1., it suffices to pick $\alpha$ and $\varepsilon$ such that $\alpha(1-\varepsilon)\geq(1+\varepsilon)$, $10<\varepsilon<$ 1, and $\alpha>1$. From these walks, a path system is constructed and given as input to 4.5 in order to construct a bramble of congestion at most $6$, and the proof of case 2. is finished.
For case 3., the goal is to apply 4.6. Let $\mathcal{I}=|V\setminus V(M_{2})|$. By assumption, $|\mathcal{I}|\geq 0.6a(a-1)$ and thus the first condition of 4.6 is satisfied. As a consequence of the applications of the Bowtie Lemma and the choice of $Z$, the authors of [19] construct, for each $(i,j)\in\mathcal{I}$, a linkage $L^{\mathfrak{Z}}_{i,j}$ such that $|L^{\mathfrak{Z}}_{i,j}|\geq d_{1}/(2^{9}\cdot 5)$ and for each pair of distinct $(i,j),(i^{\prime},j^{\prime})\in\mathcal{I}$ the graph ${\sf Int}(L^{\mathfrak{Z}}_{i,j}\cup L^{\mathfrak{Z}}f_{i^{\prime},j^{\prime}})$ is $d_{2}$-degenerate. Thus, similarly to the previous cases, to apply 4.6 we need that
@@
For reference, we now repeat the choices of $b,d_{1},d_{2}$, and $d_{3}$. We also include the choice of $a$ as it is made in [19]. Let $c$ be the constant of 4.6 and $c_{a}=c^{-4}$. In this article, we chose
@@ 4.4 Analysis of the running time
The main contributions to the running time of the algorithm come from applications of [19, Lemmas 3.5, 3.6, and 3.7] and Theorem 4.3, which depends on the LLL. The other constructions and selections done internally to the proof presented by Masařík et al. in [19, Section 4] can be done by simply following paths of the path system and picking walks and/or paths with special properties that are guaranteed to exist from the existential proofs and that can be efficiently verified. Thus, the impact of such steps on the running time is overshadowed by the contribution of the aforementioned lemmas, and we refrain from commenting on those in order to maintain a reasonable size for this article.
Let $n=|V(D)|$. As remarked above, [21, Lemmas 3.5 and 3.6] are both applied to all vertices of $G$ to build closed walks and (untangled) threaded linkages, and both naturally yield polynomial-time algorithms. The first can be realized by simply traversing the paths of each linkage $L_{i,j}$ of the $(a,b)$-path system in a particular order that guarantees that each $P\in L_{i,j}$ is traversed only once. The resulting algorithm runs in time $\mathcal{O}(a^{2}\cdot b\cdot n)$ since every $L_{i,j}$ has size $b$, $a(a-1)$ linkages are processed, and each path has size at most $n$. Similarly, the latter of those two lemmas yields an algorithm that runs in time $\mathcal{O}(b\cdot n)$ by following the greedy construction traversing the walk associated with the threaded linkage of size $b$. Thus the applications of [21, Lemmas 3.5 and 3.6] contribute to a total of $\mathcal{O}(a^{2}\cdot b\cdot n)$ to the running time of the construction.
@@
The running time of the Bowtie Lemma [19, Lemma 3.7] depends on the running time of the Partitioning Lemma of Masařík et al. [20, Lemma 11] (see also cf. [19, Lemma 3.8]). The latter consists of counting of degrees of sequences of vertices of a bipartite graph to build subgraphs with a guaranteed number of edges, and proceeding recursively on those subgraphs. The procedure yields an algorithm running in time $\mathcal{O}(\ell\cdot n^{2})$, where $\ell$ is, informally, the number of parts we want to construct. After this, the remaining part of the Bowtie Lemma is algorithmic, running in time $\mathcal{O}(b^{2}\cdot n^{2})$, since the input threaded linkages have size $b$, and thus the applications of the Bowtie Lemma contribute to a total of $\mathcal{O}(a^{4}(d_{1}\cdot n^{2}+b^{2}\cdot n^{2}))$ since it is applied to threaded linkages associated with edges of $M_{1}$ and $M_{2}$.
@@
To guarantee that the construction ends in polynomial time, the relationship $\alpha(1-\varepsilon)\geq 1+\varepsilon$ must be respected. Since $\alpha\geq 1$ is needed for the construction to be possible, this implies that $10<\varepsilon<$ 1 and $\alpha>1$. Together with the fact that $\varepsilon$ is inversely proportional to the running time stated in the 4.2, these observations imply a trade-off on the choices of $\varepsilon$ and $\alpha$. In other words, the closer $\varepsilon$ is to $1$, the faster the construction ends, and the price that we pay is increasing the dependency on $k$ that is used to find the bramble, which in turn depends on $\alpha$.
When running the algorithm, Theorem 4.3 is applied at most once since it appears once in the three cases distinguished in the proof. It is either used as part of the construction of the path system used to solve case 2., or is implicitly applied in a call to 4.6 when solving cases 1. and 3.. The running time of Theorem 4.3 is split into two parts: the construction of the input graph $G$ and then an application of 4.2. When constructing the bramble, the graph $G$ on which Theorem 4.3 is applied can be built by constructing at most $a^{4}$ intersection graphs of linkages of size $b$, and so $G$ can be built in time $\mathcal{O}(a^{4}\cdot b^{2}\cdot n)$. The dependence on $n$ is needed only to verify if two given paths intersect on a vertex and has no impact on the size of $V(G)$, which turns out to be $\mathcal{O}(a^{2}\cdot b)$ since it contains at most $a^{2}$ linkages of size $b$, and thus 4.2 runs on an input whose size depends on $k$ and $\alpha$ only.

### v2

@@ 4.3 The algorithmic construction
noticing that the running time of Theorem 4.4 (and thus of Theorems 2.14 and 2.15) is inversely proportional to $\varepsilon$, and the increase in the original dependency of $k$ needed for the construction is directly proportional to $\alpha$. Since $\alpha$ has to be at least $1$ for the choices of $b$, $x$, and $d$ above to be feasible for the construction, and $\varepsilon>0$ is needed to apply Theorem 4.4, we conclude that $10<\varepsilon<$ 1 and $\alpha>1$. Taking $\alpha$ and $\varepsilon$ satisfying those bounds and observing that $|\mathcal{I}|\geq 0.6|V|\geq 0.6a(a-1)$ (since in $V$ we excluded only the pairs $(i,i)$ with $i\in[a]$), we apply 4.7 and finish the construction.
In case 2., the goal is to apply 4.6 to a path system built in three steps. For each each $e\in M_{1}\cup M_{2}$ with endpoints $(i,j)$ and $(i^{\prime},j^{\prime})$, the Bowtie Lemma [20, Lemma 3.7] is applied to $L_{i,j}$ and $L_{i^{\prime},j^{\prime}}$ to, in short, construct sets of closed walks with bounded congestion assuming that the intersection graph of the two linkages has sufficiently large degeneracy. By the definitions of $H_{1}$ and $H_{2}$, for each $\ell\in\{1,2\}$ and $e\in M_{\ell}$ the intersection graph of the pair of threaded linkages associated with the endpoints of $e$ is not $d_{\ell}$-degenerate, each such edge is associated with a set $Z_{e}$ of closed walks of congestion at most $4$ with $|Z_{e}|\geq d_{\ell}/(2^{9}\cdot 5)$ (the denominator comes from the statement of the Bowtie Lemma). By construction of $Z$, the same holds for each $(i,j)\in Z$.
@@
If, for some $F\subseteq Z\cup M_{1}\cup M_{2}$ of size $1$ or $2$ the intersection graph ${\sf Int}(\bigcup_{g\in F}Z_{g})$ is not $d_{3}$-degenerate, then immediately apply 4.8 to construct the desired bramble and the algorithm stops. Thus, we now assume that for any such $F$ the graph ${\sf Int}(\bigcup_{g\in F}Z_{g})$is $d_{3}$-degenerate.
@@
Similarly to case 1., it suffices to pick $\alpha$ and $\varepsilon$ such that $\alpha(1-\varepsilon)\geq(1+\varepsilon)$, $10<\varepsilon<$ 1, and $\alpha>1$. From these walks, a path system is constructed and given as input to 4.6 in order to construct a bramble of congestion at most $6$, and the proof of case 2. is finished.
For case 3., the goal is to apply 4.7. Let $\mathcal{I}=|V\setminus V(M_{2})|$. By assumption, $|\mathcal{I}|\geq 0.6a(a-1)$ and thus the first condition of 4.7 is satisfied. As a consequence of the applications of the Bowtie Lemma and the choice of $Z$, the authors of [20] construct, for each $(i,j)\in\mathcal{I}$, a linkage $L^{\mathfrak{Z}}_{i,j}$ such that $|L^{\mathfrak{Z}}_{i,j}|\geq d_{1}/(2^{9}\cdot 5)$ and for each pair of distinct $(i,j),(i^{\prime},j^{\prime})\in\mathcal{I}$ the graph ${\sf Int}(L^{\mathfrak{Z}}_{i,j}\cup L^{\mathfrak{Z}}f_{i^{\prime},j^{\prime}})$ is $d_{2}$-degenerate. Thus, similarly to the previous cases, to apply 4.7 we need that
@@
For reference, we now repeat the choices of $b,d_{1},d_{2}$, and $d_{3}$. We also include the choice of $a$ as it is made in [20]. Let $c$ be the constant of 4.7 and $c_{a}=c^{-4}$. In this article, we chose
@@ 4.4 Analysis of the running time
The main contributions to the running time of the algorithm come from applications of [20, Lemmas 3.5, 3.6, and 3.7] and Theorem 4.4, which depends on the LLL. The other constructions and selections done internally to the proof presented by Masařík et al. in [20, Section 4] can be done by simply following paths of the path system and picking walks and/or paths with special properties that are guaranteed to exist from the existential proofs and that can be efficiently verified. Thus, the impact of such steps on the running time is overshadowed by the contribution of the aforementioned lemmas, and we refrain from commenting on those in order to maintain a reasonable size for this article.
Let $n=|V(D)|$. As remarked above, [22, Lemmas 3.5 and 3.6] are both applied to all vertices of $G$ to build closed walks and (untangled) threaded linkages, and both naturally yield polynomial-time algorithms. The first can be realized by simply traversing the paths of each linkage $L_{i,j}$ of the $(a,b)$-path system in a particular order that guarantees that each $P\in L_{i,j}$ is traversed only once. The resulting algorithm runs in time $\mathcal{O}(a^{2}\cdot b\cdot n)$ since every $L_{i,j}$ has size $b$, $a(a-1)$ linkages are processed, and each path has size at most $n$. Similarly, the latter of those two lemmas yields an algorithm that runs in time $\mathcal{O}(b\cdot n)$ by following the greedy construction traversing the walk associated with the threaded linkage of size $b$. Thus the applications of [22, Lemmas 3.5 and 3.6] contribute to a total of $\mathcal{O}(a^{2}\cdot b\cdot n)$ to the running time of the construction.
@@
The running time of the Bowtie Lemma [20, Lemma 3.7] depends on the running time of the Partitioning Lemma of Masařík et al. [21, Lemma 11] (see also cf. [20, Lemma 3.8]). The latter consists of counting of degrees of sequences of vertices of a bipartite graph to build subgraphs with a guaranteed number of edges, and proceeding recursively on those subgraphs. The procedure yields an algorithm running in time $\mathcal{O}(\ell\cdot n^{2})$, where $\ell$ is, informally, the number of parts we want to construct. After this, the remaining part of the Bowtie Lemma is algorithmic, running in time $\mathcal{O}(b^{2}\cdot n^{2})$, since the input threaded linkages have size $b$, and thus the applications of the Bowtie Lemma contribute to a total of $\mathcal{O}(a^{4}(d_{1}\cdot n^{2}+b^{2}\cdot n^{2}))$ since it is applied to threaded linkages associated with edges of $M_{1}$ and $M_{2}$.
@@
For $\alpha=1$ we can apply 4.2 to obtain Theorem 2.16.
To guarantee that the construction ends in polynomial time when using Theorem 4.4, the relationship $\alpha(1-\varepsilon)\geq 1+\varepsilon$ must be respected. Since $\alpha\geq 1$ is needed for the construction to be possible, this implies that $10<\varepsilon<$ 1 and $\alpha>1$. Together with the fact that $\varepsilon$ is inversely proportional to the running time stated in the 4.3, these observations imply a trade-off on the choices of $\varepsilon$ and $\alpha$. In other words, the closer $\varepsilon$ is to $1$, the faster the construction ends, and the price that we pay is increasing the dependency on $k$ that is used to find the bramble, which in turn depends on $\alpha$.
When running the algorithm, Theorem 4.4 is applied at most once since it appears once in the three cases distinguished in the proof. It is either used as part of the construction of the path system used to solve case 2., or is implicitly applied in a call to 4.7 when solving cases 1. and 3.. The running time of Theorem 4.4 is split into two parts: the construction of the input graph $G$ and then an application of 4.3. When constructing the bramble, the graph $G$ on which Theorem 4.4 is applied can be built by constructing at most $a^{4}$ intersection graphs of linkages of size $b$, and so $G$ can be built in time $\mathcal{O}(a^{4}\cdot b^{2}\cdot n)$. The dependence on $n$ is needed only to verify if two given paths intersect on a vertex and has no impact on the size of $V(G)$, which turns out to be $\mathcal{O}(a^{2}\cdot b)$ since it contains at most $a^{2}$ linkages of size $b$, and thus 4.3 runs on an input whose size depends on $k$ and $\alpha$ only.